Butane, C4H10, is used as a fuel for barbecues and as an aerosol propellant. If you have 108mL of butane at 715 mmHg and 25oC, what is the mass, in grams, of butane? Show all work!
Answer #1

given:

P = 715 mmHg

= (715/760) atm

P = 0.9407 atm

T = 25 + 273 K = 298 K

V = 108 mL = 0.108 L , R = 0.0821

and

n = PV/RT
= 0.9407*0.108/0.0821*298

n = 0.004155 moles

now,

mass of butane = molar mass of butane*moles

= 58*0.004155

mass of butane = 0.241 gm ............. answer